OpenStudy (user123):

when solving parametric equations, exactly when do we need to eliminate the t?

OpenStudy (user123):

do we need to eliminate over here?

zepdrix (zepdrix):

I don't think you ever "need to" eliminate the parameter. You can simply choose to do so when it's convenient. In this case, we have a squared t in each equation, it might be fairly difficult to eliminate the parameter.

zepdrix (zepdrix):

\[\large\rm \frac{dy}{dx}=\frac{(dy/dt)}{(dx/dt)}\]So in this problem, it's probably easier to just take your derivatives separately, and then plug them in.\[\large\rm y=t^2-t+1\qquad\to\qquad \frac{dy}{dt}=2t-1\]So that's what would go in our numerator, ya?\[\large\rm \frac{dy}{dx}=\frac{2t-1}{(dx/dt)}\]

OpenStudy (user123):

yes. \[\frac{ 2t-1 }{ 2t }\]

OpenStudy (user123):

we're not allowed to eliminate 2t together, right? what would be the next step?

zepdrix (zepdrix):

Woops, your denominator doesn't look quite right.

zepdrix (zepdrix):

You got 0 for the derivative of 1/t? O_o whuuuu

OpenStudy (user123):

\[\frac{ 0 \times v - 1 \times 1 }{ t^2} = \left(\begin{matrix}-1 \\ t^2\end{matrix}\right)\]

zepdrix (zepdrix):

\[\large\rm \frac{dy}{dx}=\frac{2t-1}{-\frac1{t^2}+2t}\]Mmm ok that looks better.

OpenStudy (user123):

\[\frac{ 2t-1 }{ \frac{ -1 }{ t^2 } }\]

OpenStudy (user123):

where'd you get the 2t?

zepdrix (zepdrix):

Where did you get the 2t from the first time?

OpenStudy (user123):

Ohhh I see.

OpenStudy (user123):

It's from the original equation I'm sorry

zepdrix (zepdrix):

So we have a derivative function.. that we need to evaluate at (x,y)=(2,1). This next step might seem a little confusing at first. We don't have any x or y in our derivative function! So what can we do? Well, we can plug our (x,y) into our original system to find the corresponding t value.

zepdrix (zepdrix):

\[\large\rm x=\frac1t+t^2\qquad\qquad\qquad y=t^2-t+1\]Plugging in (2,1) gives us,\[\large\rm 2=\frac1t+t^2\qquad\qquad\qquad 1=t^2-t+1\]

OpenStudy (user123):

Oh? But that still does not make sense to me

zepdrix (zepdrix):

Well let's look at the second equation. Maybe we can come up with some values for t.\[\large\rm 1=t^2-t+1\]Subtracting 1 from each side,\[\large\rm 0=t^2-t\]From here, maybe factor a t out of each term,\[\large\rm 0=t(t-1)\]

zepdrix (zepdrix):

Do you see the t values? There should be two of them.

OpenStudy (user123):

1 and 0

OpenStudy (user123):

\[t+t^2=2\]

OpenStudy (user123):

t = 1

OpenStudy (user123):

Only 1? Does it have to satisfy both equations?

zepdrix (zepdrix):

\[\large\rm 2=\frac1t+t^2\]Well it turns out that t=0 does not satisfy this other equation. Plugging in t=0 would give us this undefined value 1/0. So that's great! It gives us our one and only value for t, t=1. Yes, it has to satisfy both equations.

zepdrix (zepdrix):

So we've determined that the point (x,y) = (2,1) is the time when t=1. So that's when we'll evaluate our derivative function, at t=1.

OpenStudy (user123):

The entire answer to this problem seems to be 1.

OpenStudy (user123):

I plugged in 1 for dy/dx and got 1.

zepdrix (zepdrix):

\[\large\rm \frac{dy}{dx}=\frac{2t-1}{-\frac1{t^2}+2t}\]Umm let's see..\[\large\rm \frac{dy}{dx}=\frac{2(1)-1}{-\frac1{1^2}+2(1)}=\frac11\]Mmmm yes good job!

OpenStudy (user123):

Thank you for helping :)

zepdrix (zepdrix):

np